Find the equation of the line which satisfies each set of conditions:
passes through the point with coordinates (-2, 1) and is parallel to the line, 3y - x = 9

Answer by plourde216(7)

In order to complete this problem, you must first solve the equation3y-x=9 for y to get it into slope-intercept form.
3y-x=9
3y=9+x
y=3+(1/3)x
Now you are able to determine the slope of this line. The slope is always the number in front of the x. In this case it is (1/3). Next, you must determine the relationship between the slopes of parallel lines. All parallel lines will have slopes that are equal. Therefore we know that the equation of the line we must find will have a slope of (1/3). Because we know the slope and a point on the line, we can plug our information into the point-slope formula.
y-1=(1/3)(x+2)
y-1=(1/3)x+(2/3)
y=(1/3)x + (5/3)
I hope that this helps